# Inverse-Squeezing Kennedy Receiver for Near-Helstrom Discrimination of Displaced-Squeezed BPSK

### v1

@@ V IS-Kennedy Receiver with Squeezing Mismatch
VI Conclusion
In this paper, we proposed and analyzed an Inverse-squeezing Kennedy (IS-Kennedy) receiver for the discrimination of S-BPSK signals. By introducing an inverse-squeezing module and photon-number-resolving (PNR) detection technology, this scheme not only significantly outperforms conventional coherent-state receivers in the low-energy regime but also approaches the quantum limit (Helstrom Bound) of S-BPSK under ideal conditions. Addressing non-ideal factors in practical applications, we established a comprehensive theoretical model incorporating detection efficiency, dark counts, and squeezing mismatch. Our research reveals that the PNR detector effectively mitigates performance degradation caused by dark counts by adaptively adjusting the decision threshold. Furthermore, under squeezing mismatch conditions, a high-resolution PNR detector leverages the parity photon statistics of the squeezed vacuum state to reduce the saturation error probability by several orders of magnitude. These results demonstrate that the IS-Kennedy receiver offers a highly promising implementation scheme for high-sensitivity quantum communication based on squeezed states.
In this paper, we proposed an inverse-squeezing Kennedy (IS-Kennedy) receiver for discriminating phase-opposite displaced squeezed-vacuum states (S-BPSK). By applying a Kennedy-type nulling displacement followed by a matched inverse-squeezing operation (the inverse of the transmitter squeezing), the receiver converts the squeezing resource into an effective separation gain between hypotheses, enabling photon counting to extract the information carried jointly by displacement and squeezing. With a photon-number-resolving (PNR) detector of resolution M and a maximum a posteriori (MAP) decision rule based on the count statistics, the IS-Kennedy receiver can closely approach the S-BPSK Helstrom limit under ideal conditions and substantially outperform coherent-state baselines in the low-energy regime.
We further developed a unified performance analysis that incorporates practical imperfections, including finite detection efficiency, dark counts, finite PNR resolution, and inverse-squeezing mismatch (e.g., errors in the squeezing parameter and angle). The results reveal two key behaviors. First, in the presence of dark counts, PNR-based MAP detection improves robustness through an adaptive integer threshold, leading to characteristic step-like performance as system parameters vary. Second, inverse-squeezing mismatch introduces a parity-structured saturation (error floor) that is intrinsic to the photon-number statistics of squeezed states; importantly, increasing the PNR resolution suppresses this mismatch-induced floor in “parity steps,” with pronounced gains when the resolution crosses even-photon thresholds. Together, these findings indicate that the IS-Kennedy architecture provides a practical path toward near-optimal squeezed-state receivers using currently accessible operations—displacement, (inverse) squeezing, and photon counting.
Finally, while our model captures the dominant non-idealities for photon-counting implementations, additional experimental effects—such as displacement calibration error, local-oscillator phase noise, temporal/spectral mode mismatch, finite detector dead time, and afterpulsing—remain to be incorporated. Addressing these effects and exploring adaptive/feedback implementations constitute promising directions for future work.
Appendix A The Wigner Function of Displaced Squeezed Vacuum States
The Displaced Squeezed States (DSSs) discussed in this paper belong to single-mode Gaussian states, whose Wigner function is a bivariate Gaussian distribution fully characterized by two parameters: the displacement vector $\vec{d}$ and the covariance matrix $\mathbf{V}$. Given a DSS denoted as $\left|\psi\right\rangle$, its quantum mechanical expression is
$\left|\psi\right\rangle=\left|\alpha,r\right\rangle=D(\alpha)S(r)\left|0\right\rangle,\qquad\alpha,r\in\mathbb{R}_{+}.$ (50)
The first parameter, the displacement vector $\vec{d}$, is a two-dimensional vector representing the translation in phase space relative to the origin. Its components correspond to the expectation values of the quadrature operators $\hat{X},\hat{P}$:
$\vec{d}=\begin{pmatrix}d_{x}\\ d_{p}\end{pmatrix}=\begin{pmatrix}\left\langle\hat{X}\right\rangle\\[2.0pt] \left\langle\hat{P}\right\rangle\end{pmatrix}$. (51)
For $\left|\psi\right\rangle$, $d_{x}=\sqrt{2}\,\text{Re}(\alpha)=\sqrt{2}\alpha$ and $d_{p}=\sqrt{2}\,\text{Im}(\alpha)=0$. The second parameter, the covariance matrix $\mathbf{V}$, is a $2\times 2$ real symmetric positive-definite matrix describing the “shape” and “size” of the quantum state in phase space. Its elements are variances and covariances:
$\mathbf{V}$ $=\begin{bmatrix}V_{xx}&V_{xp}\\ V_{px}&V_{pp}\end{bmatrix}$ (52)
$=\begin{bmatrix}\left\langle(\Delta\hat{X})^{2}\right\rangle&\frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle\\[4.0pt] \frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle&\left\langle(\Delta\hat{P})^{2}\right\rangle\end{bmatrix}$,
where $\Delta\hat{O}=\hat{O}-\left\langle\hat{O}\right\rangle$, and the anticommutator is defined as $\left\{\hat{A},\hat{B}\right\}=\hat{A}\hat{B}+\hat{B}\hat{A}$. For $\left|\psi\right\rangle$, the covariance matrix is
$\mathbf{V}=\begin{bmatrix}\frac{1}{2}e^{-2r}&0\\[2.0pt] 0&\frac{1}{2}e^{2r}\end{bmatrix}$, (53)
indicating that in phase space, the state is an ellipse (coherent states and vacuum states appear as circles), with fluctuations squeezed along the $X$-direction and expanded along the $P$-direction, while the overall phase-space area remains unchanged ($\det\mathbf{V}=\tfrac{1}{2}e^{-2r}\cdot\tfrac{1}{2}e^{2r}=\tfrac{1}{4}$).
Substituting $\vec{d}$ and $\mathbf{V}$ into the general formula yields the Wigner function for $\left|\psi\right\rangle$:
$\begin{aligned} &W(\vec{r}|\psi)=\frac{1}{2\pi\sqrt{\det\mathbf{V}}}\,\exp\left\{-\frac{1}{2}(\vec{r}-\vec{d})^{T}\mathbf{V}^{-1}(\vec{r}-\vec{d})\right\}\\ &\to W(x,p|\psi)=\frac{1}{\pi}\exp\left\{-\,\text{e}^{2r}\big(x-\sqrt{2}\alpha\big)^{2}-\,\text{e}^{-2r}p^{2}\right\},\end{aligned}$ (54)
where $\vec{r}=\begin{pmatrix}x\\ p\end{pmatrix}$ is a point in phase space, and $\det\mathbf{V}$ denotes the determinant of $\mathbf{V}$.
@@ Appendix A The Wigner Function of Displaced Squeezed Vacuum States
Appendix B Fock-basis Population of Squeezed Vacuum and Displaced Squeezed States
@@ Acknowledgements.
This work was supported by Independent Innovation Science Fund Program of National University of Defense Technology, China (Grant No. 22-ZZCX-036), Key research & development program of Guangxi (Grant No. GuiKeAB23075155), and the Key Research & Development Program of Guangxi (Grant No. AB23075112).
@@
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.

### v2

@@ II.2 Receiver under squeezing mismatch condition
III Simulation methods and results
The numerical results were obtained using QuTiP in Python [6]. We model the receiver output states in a truncated Fock space and compute the photon-number distributions from the corresponding density operators. Unless otherwise specified, the simulations use a cutoff $N_{\text{cut}}=100$. We verified numerical convergence by increasing the cutoff dimension until the resulting error probabilities changed negligibly. For each mean photon number $N$, the S-BPSK states were generated with the optimal squeezing fraction $\beta_{\text{opt}}=\frac{N}{2N+1}$.
Figure 2: Performance analysis of the IS-Kennedy receiver under inverse-squeezing mismatch conditions. (a) Error probability of IS-Kennedy with SPD. (b) Error probability of IS-Kennedy with PNR(3). (c) The performance ratio (in dB) of IS-Kennedy with PNR(10) relative to the S-BPSK SQL. (d) The Fock-basis population of $\zeta_{0}^{\text{sm}},\ \zeta_{1}^{\text{sm}}$ with inverse-squeezing mismatch of $\Delta r=0.02,\ \Delta\theta=0.03\pi$.
Figure 2(a) and Fig. 2(b) show the error probability under SPD and PNR(3), respectively. In the ideal case, the error decreases rapidly with $N$, reflecting the benefit of the inverse-squeezing transformation. Under amplitude-only mismatch, $\left(\Delta r,\Delta\theta\right)=\left(0.02,0\right)$, the performance remains close to the ideal curve at low $N$ but saturates at high $N$, as expected from the residual squeezing acting on the nulled state $\zeta_{0}^{\text{sm}}$. The saturation levels agree with Eq. (30): for SPD the error floor is about $10^{-4}$, whereas for PNR(3) it is reduced to about $3\times 10^{-8}$. This confirms the parity-step behavior predicted by the small-mismatch analysis, namely that increasing the PNR resolution suppresses the amplitude-mismatch-induced error only when the resolution crosses an even-photon threshold. The physical origin is illustrated in Fig. 2(d): under mismatch, $\zeta_{0}^{\text{sm}}$ becomes a squeezed vacuum with only even-photon components. As a result, the likelihood function is generally non-monotonic in photon number, and the MAP decision is no longer a simple single-threshold rule.
Phase mismatch causes a much stronger degradation. For $\left(\Delta r,\Delta\theta\right)=\left(0,0.03\pi\right)$, the error begins to deviate from the ideal case at moderate $N$ and increases significantly at higher energies. Moreover, the curves for $\left(0,0.03\pi\right)$ and $\left(0.02,0.03\pi\right)$ nearly overlap in both Fig. 2(a) and Fig. 2(b), showing that once phase mismatch is present, the additional effect of amplitude mismatch is comparatively weak. This is consistent with Eq. (19), where the phase-mismatch contribution is amplified by $\sinh\left(2r\right)/2$ and therefore dominates the residual squeezing strength $r_{m}$ at sufficiently large squeezing (i.e. at high $N$). Although finite-resolution PNR clearly outperforms SPD in the mismatched regime, it does not remove the dominant sensitivity to phase mismatch.
Figure 2(c) further plots the error-probability ratio in dB relative to the S-BPSK SQL, where negative values indicate performance below the SQL. In the ideal case, the IS-Kennedy receiver preserves a clear advantage as $N$ increases. Under amplitude-only mismatch, this advantage is largely maintained over a broad range of $N$, but is eventually limited by the saturation effect. By contrast, phase mismatch rapidly reduces the SQL advantage, and the error probability even continues to rise at large $N$.
Figure 3: The maximum tolerable inverse-squeezing mismatch for which the IS-Kennedy receiver still outperforms the S-BPSK SQL, i.e., $P_{\text{err}}^{\text{ISK}}<P_{\text{SQL}}^{\text{S-BPSK}}$. (a) Maximum tolerable amplitude mismatch $\Delta r_{\text{max}}$ versus mean photon number $N$. (b) Maximum tolerable phase mismatch $\Delta\theta_{\text{max}}$ versus NN. Results are shown for PNR(1), PNR(2), and PNR(3).
Figure 3 provides a more direct robustness metric by showing the maximum tolerable inverse-squeezing mismatch for which the IS-Kennedy receiver still outperforms the S-BPSK SQL, i.e., $P_{\text{err}}^{\text{ISK}}<P_{\text{SQL}}^{\text{S-BPSK}}$. As shown in Fig. 3(a), the tolerable amplitude mismatch $\Delta r_{\text{max}}$ is non-monotonic in $N$: it first increases, reaches its largest value in the low-to-intermediate energy region, and then decreases at higher $N$. For PNR(1) and PNR(2), the admissible range rapidly shrinks at large $N$, whereas PNR(3) retains an additional tolerance window. A similar trend is observed for the phase mismatch in Fig. 3(b), but the tolerable phase error collapses more rapidly with increasing $N$, in agreement with the enhanced phase sensitivity predicted by Eq. (19). Moreover, the extra tolerance window preserved by PNR(3) is consistent with the parity-step behavior discussed above: increasing the PNR resolution improves mismatch robustness in a non-smooth, threshold-like manner rather than continuously.
Overall, the simulations show that the IS-Kennedy receiver is far more sensitive to phase mismatch than to amplitude mismatch, and that finite-resolution PNR improves robustness mainly against amplitude mismatch rather than imperfect phase locking.
IV Conclusion
In conclusion, we have investigated inverse-squeezing mismatch in the IS-Kennedy receiver for the discrimination of binary displaced squeezed vacuum states. We showed that imperfect inverse squeezing is equivalent to a residual squeezing after nulling, which modifies the output photon-number statistics and makes the optimal MAP decision generally non-single-threshold. The analysis further reveals a strong asymmetry between the two mismatch types: the receiver is much more sensitive to phase mismatch than to amplitude mismatch, because the phase error is effectively amplified by the squeezing process. Numerical results confirm that amplitude-only mismatch leads to a saturation error floor, and that finite-resolution PNR detection suppresses this floor in a parity-step manner rather than continuously with detector resolution. By contrast, phase mismatch rapidly degrades the receiver performance and can destroy its advantage over the S-BPSK SQL at high signal energies. The maximum-tolerable-mismatch results further show that the practically useful operating region of the IS-Kennedy receiver is ultimately limited by phase locking. These results clarify the physical origin of mismatch-induced degradation and provide guidance for the experimental design of inverse-squeezing receivers with finite-resolution photon counting.
@@ IV Conclusion
Appendix A Fock-basis Population of Squeezed Vacuum and Displaced Squeezed States
@@ Acknowledgements.
This work was supported by Independent Innovation Science Fund Program of National University of Defense Technology, China (Grant No. 22-ZZCX-036), Key research & development program of Guangxi (Grant No. GuiKeAB23075155), Innovative Talent Development Fund of Information Support Force Engineering University (Grant No. XJKT-QT-25-02-GW), and the Key Research & Development Program of Guangxi (Grant No. AB23075112).
@@
The data that support the findings of this study are available from the authors upon request.
Code Availability
The numerical simulations in this study were performed in Python using QuTiP 5.1.1. The custom scripts used to generate the results are available from the corresponding author on reasonable request.
Competing interests
The authors declare no competing interests.